# Ancestor regression in structural vector autoregressive models

### v1

@@ A.5.1 Proof of Lemma 1
$\mathbf{e}_{k}^{\top}\sum_{\tau=m+1}^{\infty}\tilde{\mathbf{B}}^{\tau}\mathbb{E}\mathopen{}\mathclose{{}\left[\boldsymbol{\xi}_{t\_p}\boldsymbol{\xi}_{t\_p}^{\top}}\right]\mathopen{}\mathclose{{}\left(\tilde{\mathbf{B}}^{\tau}}\right)^{\top}\mathbf{e}_{k}$ $\leq\sum_{\tau=m+1}^{\infty}\mathopen{}\mathclose{{}\left\|\mathbf{e}_{k}^{\top}\tilde{\mathbf{B}}^{\tau}}\right\|_{2}^{2}\lambda_{\max}\mathopen{}\mathclose{{}\left(\mathbb{E}\mathopen{}\mathclose{{}\left[\boldsymbol{\xi}_{t\_p}\boldsymbol{\xi}_{t\_p}^{\top}}\right]}\right)$
$\leq C\sum_{\tau=m+1}^{\infty}\lambda_{\max}\mathopen{}\mathclose{{}\left(\tilde{\mathbf{B}}}\right)^{2\tau},$
@@ A.7 Combined p-values
Appendix B Details on the simulation setup
We use the following distributions for the $\epsilon_{t,j}$: two $t_{7}$ distributions, two centered uniform distributions, a centered Laplace distribution with scale $1$, and a standard normal distribution. All distributions are normalized to have unit variance. For each simulation run, we randomly permute the distributions to assign them to $\epsilon_{t,1}$ to $\epsilon_{t,6}$.

### v2

@@ A.5.1 Proof of Lemma 1
$=\mathbf{e}_{k}^{\top}\sum_{\tau=m+1}^{\infty}\tilde{\mathbf{B}}^{\tau}\mathbb{E}\mathopen{}\mathclose{{}\left[\boldsymbol{\xi}_{t\_p}\boldsymbol{\xi}_{t\_p}^{\top}}\right]\mathopen{}\mathclose{{}\left(\tilde{\mathbf{B}}^{\tau}}\right)^{\top}\mathbf{e}_{k}$
$\leq\sum_{\tau=m+1}^{\infty}\mathopen{}\mathclose{{}\left\|\mathbf{e}_{k}^{\top}\tilde{\mathbf{B}}^{\tau}}\right\|_{2}^{2}\lambda_{\max}\mathopen{}\mathclose{{}\left(\mathbb{E}\mathopen{}\mathclose{{}\left[\boldsymbol{\xi}_{t\_p}\boldsymbol{\xi}_{t\_p}^{\top}}\right]}\right)\leq C\sum_{\tau=m+1}^{\infty}\lambda_{\max}\mathopen{}\mathclose{{}\left(\tilde{\mathbf{B}}}\right)^{2\tau},$
@@ A.7 Combined p-values
Appendix B Simulation results
B.1 Details on the simulation setup
We use the following distributions for the $\epsilon_{t,j}$: two $t_{7}$ distributions, two centered uniform distributions, a centered Laplace distribution with scale $1$, and a standard normal distribution. All distributions are normalized to have unit variance. For each simulation run with $6$ variables in Sections 2.1 and 3.3, we randomly permute the distributions to assign them to $\epsilon_{t,1}$ to $\epsilon_{t,6}$. For the simulation runs with more variables in Appendix B.2, we randomly draw distributions with replacement from those $6$.
@@ B.1 Details on the simulation setup
For the simulations with a hidden variable, we randomly sample a variable and exclude its time series from the data to estimate the models. This hidden variable is not counted to calculate the power. However, to obtain the true ancestors, it is considered, i.e., the ancestors of the unobserved variable are true ancestors of the unobserved variable’s descendants.
B.2 Simulations with more time series
We repeat the simulations in Section 3.3 with more time series, namely $d=10$ and $d=50$.
In Figure 10, we show the performance when all time series are observed. For $d=50$, we cannot do estimation with only $100$ observations, so this is omitted in the simulation study. As we randomly sample from the distributions as outlined above, there is the possibility that several Gaussian distributions are assigned in a simulation run, and for $d=50$ this is very likely. Hence, LiNGAM has no consistency guarantee anymore, i.e., the error rate can remain high even for larger sample sizes, and ancestor regression may lose power to detect instantaneous effects. The Bonferroni-Holm correction together with the correction to obtain the summary p-values leads to a very high multiplier for $d=50$. Hence, the p-values for non-ancestors, which are approximately uniformly distributed, are mostly $1$ after correction, and the error rate is always low. Furthermore, there are not many true nulls for the summary effects which could lead to type I errors.
(a) $d=10$
(b) $d=50$
Figure 10: Nodewise ancestor detection in a structural vector autoregressive model of order $p=1$ with $d$ variables. See the caption of Figure 4 for explanations.
The results when setting one time series at random as unobserved are shown in Figure 11. As for $d=6$, we lose error control when one time series is unobserved. There is still some separation between ancestors and non-ancestors, i.e., our test statistic provides an indication of what could be true ancestors. However, for $d=50$, it requires a very high sample size for reasonable performance. As outlined before, the potential for type I errors is lower for the summary effects. This applies even with an unobserved time series.
(a) $d=10$
(b) $d=50$
Figure 11: Nodewise ancestor detection in a structural vector autoregressive model of order $p=1$ with $d$ variables, whereof one is unobserved. See the caption of Figure 4 for explanations.